OpenStudy (anonymous):

Through how many radians does a clock turn in 48 minutes

OpenStudy (unklerhaukus):

which hand of the clock are we looking at?

OpenStudy (unklerhaukus):

all the way around = 2π[radians] = 360[°] = 60 [min]

OpenStudy (anonymous):

its not asking which hand

OpenStudy (anonymous):

so 60 min in a hour and the circle measures 2pi radians

OpenStudy (unklerhaukus):

in 48 minutes the minute hand will have gone forty-eight sixtieths of the way around ie 48[min] = (48/60)*60 [min] = (48/60)*360[°] = (48/60)*2π [radians]

OpenStudy (anonymous):

have another question for you solve l 3x-2 l < 4 will the outcome end up x<2

OpenStudy (unklerhaukus):

thats only part of the solution